OpenStudy (anonymous):

y=-cos(pi/6) I know the answer is -sqt3/2 but how do you get that answer?

OpenStudy (anonymous):

i know i can put it into ym calculator, but the answer isnt in pi form. Im not sure how to do so

OpenStudy (anonymous):

This kind of follows from what the cosine function is defined as. It happens that the \(x\) coordinate of a point on the unit circle that intersects with a line that has an angle of \(\frac{\pi}{6}\) with the x axis is \(\sqrt{\frac{3}{2}}\)

OpenStudy (anonymous):

i think i get it now:)

OpenStudy (anonymous):

Awesome.

OpenStudy (anonymous):

how do you determine the domain?

OpenStudy (anonymous):

pi is equal to 180 degrees. 2 pi is a complete circle which is 360 degrees. so you should understand that pi = 180 degrees. y = -cos(180/6) so we are talking about a 30 degree angle. y = - cos(30 degree) angle is the same as cos 30 degree angle. There are some basic trig functions you must memorize. One of these is cos (pi/6) = sqrt(3)/2 I don't think the answer is -sqrt(3)/2 I think it is sqrt(3)/2 Cosine has the same sign in the 1st and 4th quadrants which is where this is

OpenStudy (anonymous):

The domain of a function is the \(x\) values over which it is defined. The \(\cos\) function happens to be defined over all \(x\), so the domain would be \(-\infty<x<\infty\).